# AAAI Press Formatting Instructions for Authors Using — A Guide

### v2

@@ Abstract
Clinical decision-making often involves interpreting images (e.g., radiology) for making diagnoses. Retrieving relevant visual information from medical literature and hospital records could enhance diagnostic accuracy. In this paper, we develop a model in which a multimodal retriever is jointly optimized with an LVLM for medical diagnosis, unlike standard RAG where LVLM error signal is not propagated down to the retriever. We show that using only general-purpose backbones, with only lightweight fine-tuning, our model is able to achieve competitive results with medically-pretrained models across clinical multi-label classification and visual question answering tasks. In a novel analysis, we additionally find that in many cases different top retrieved images each lead to different predictions for a given target, and that these cases are empirically challenging for all models, even for non-retrieval models. Our joint retrieval optimization significantly improves these challenging cases over standard RAG. However, oracle analysis reveals that while the correct diagnosis is frequently achievable using one of the top retrieved images, in practice there is a large performance gap from the oracle, and rerankers using frontier LVLMs do not close this gap–leaving ample room for improvement by future methods.Code available at: https://github.com/Nirmaz/JOMED.
Introduction
@@
Figure 1: We jointly optimize a multimodal retriever and a Large Vision-Language Model (LVLM) for medical tasks. Our model achieves competitive results without resource-intensive medical pre-training and significantly improves performance on challenging cases where different retrieved images lead to inconsistent retrieval prediction.
@@
Inferring diagnoses from medical imagery is a fundamental part of clinical decision-making. Large Vision Language Models (LVLMs) have been widely explored for medical diagnosis (Thawakar et al. 2024; Wu et al. 2023; Zhang et al. 2023b; Moor et al. 2023; Li et al. 2023). To improve the performance of LVLMs in the medical field, retrieval augmentation has been adopted and has shown promising results, providing more accurate and also explainable methods (He et al. 2024a; Xia et al. 2024b, a; Wu et al. 2025).
@@
In this work, we explore a lightweight fine-tuning approach for taking a general-purpose LVLM and optimizing it jointly with a general-purpose multimodal retriever. Different from the common retrieval augmented generation (RAG) setup, we perform joint optimization of the LVLM and retriever. This joint optimization teaches the LVLM to utilize retrieved knowledge for making accurate medical predictions while at the same time tuning the multimodal retriever to retrieve information that leads to correct predictions by the LVLM. The retriever model retrieves images as well as captions and clinical reports from both literature and hospital records. Our method involves sequential multimodal training with a dual-head retriever architecture and a retrieval loss that enhances the retriever’s ability to provide more predictive context to the LVLM.
Poorly optimized retrieval mechanisms can mislead models (Yoran et al. 2023; Sun et al. 2024). Figure 2 demonstrates this problem with an ultrasound image that shows no breast cancer. When the retrieval system selects an irrelevant reference image showing cancer features, the model incorrectly identifies cancer in the input image. When we jointly optimize the retriever to find more relevant images and text, the model’s prediction flips to the correct diagnosis.
Using our approach, we are able to achieve competitive results in medical image classification and visual question answering (VQA). Importantly, our focus in this work is not to chase SOTA results but to shed light on several interesting observations. One, that a simple and effective joint LVLM-retriever optimization mechanism can lead to a substantial boost in results over standard multimodal RAG and should thus be more widely adopted for medical diagnosis.
@@
Second, that this mechanism achieves competitive results by using models with no medical pre-training (for neither the LVLM nor the retriever) and only lightweight fine-tuning; this resonates with recent findings showing general-purpose LVLMs can rival their medical counterparts (Jeong et al. 2024) and has potentially important implications considering the resource-intensive nature of pre-training processes.
@@
Third, previous work on joint training in the general domain required large-scale pre-training followed by few-shot task-specific adaptation (Izacard et al. 2023; Hu et al. 2023; Lin et al. 2024). In contrast, our novelty is in conducting joint training directly on downstream tasks with no pretraining and only lightweight fine-tuning (as few as 546 samples). This provides first evidence about the feasibility and utility of lightweight data-efficient joint training directly on downstream tasks as opposed to in the pre-training setting. We are also the first to explore and show the utility of joint training in the multimodal medical domain.
@@
Fourth and importantly, we conduct a novel analysis that finds that our method helps in particular to address a class of errors we term inconsistent retrieval predictions. Inconsistent retrieval predictions are cases in which for a given patient image, each retrieved image leads the model to make different predictions. These cases commonly occur across our experiments and are substantially more difficult for models, both retrieval-augmented and non-retrieval models. This instability with respect to different retrieval candidates leads to high prediction entropy/uncertainty across classes, degrading not only overall RAG results but also their reliability (Lambert et al. 2022).
As we show, our joint training mechanism significantly mitigates this issue and achieves large improvement over standard fine-tuned RAG on these cases.
We further discover that for a large proportion of these inconsistent cases, at least one retrieved candidate enables the model to predict the correct answer in an oracle setting in which we provide the model with the correct retrieved candidate. This suggests useful information is being retrieved but is often lost among less helpful candidates. To address this, we explored using the powerful multimodal model o3 (OpenAI 2025) as a reranker to identify the most predictive instances. We found it did not close the performance gap to the oracle and was comparable or inferior to our method, leaving significant room for future improvement.
@@
  • We develop a method where a multimodal retriever and a Large Vision-Language Model (LVLM) are jointly optimized for medical classification and VQA. Unlike standard Retrieval-Augmented Generation (RAG), our method propagates error signal from the LVLM down to the retriever while training on downstream tasks.
@@
  • We show that our method achieves competitive results compared to medically pre-trained models by using only general-purpose backbones. Unlike previous general-domain joint training methods that require pre-training, we conduct joint training directly on downstream tasks with only lightweight data-efficient fine-tuning.
@@
  • In a novel analysis we identify inconsistent retrieval predictions, challenging cases where different retrieved images lead to different predictions for the same query. Our joint optimization method significantly improves performance on these cases when compared to standard RAG.
@@
Figure 2: Joint training impacts inconsistent retrieval predictions. After joint training, retrieved images show greater alignment with query image labels. In Breast, a cancer-free ultrasound query (a) initially retrieved a lesion image (b), however, after joint training, the retrieved image is less directly related to the wrong label (c). For Retina, joint training changed retrieval from a different modality (e) to be more similar to the query (d). In VQA-RAD, retrieval shifted from an unrelated medical condition (h) to an image depicting the condition of the query image.
Related Work
@@
Retrieval Augmented Joint Training has been explored primarily in unimodal (text only) NLP with work such as ATLAS (Izacard et al. 2023), which conducted large-scale pretraining of both reader and retriever models and evaluated in zero-shot and few-shot scenarios in the general domain. REVEAL (Hu et al. 2023) extended this to the multimodal setting, achieving state-of-the-art results in visual tasks in the general domain with extensive pre-training. (Shi et al. 2023) proposed training textual retrievers based on reader performance, influencing subsequent joint-training methods (Lin et al. 2024). introduced another joint training variant and proposed sequential training: first training the reader with retrieval augmentation, then training the retriever. Siriwardhana et al. (2023) demonstrated joint training’s effectiveness for domain adaptation. Our work differs by being the first to explore joint training optimized directly for downstream tasks, with minimal fine-tuning, as opposed to requiring large-scale pre-training. Our extensive novel analyses are the first to shed light on inconsistent predictions and how joint training on downstream tasks helps address them.
@@
Multimodal Retrieval Augmentation in Medical Applications. Multimodal retrieval augmentation began with encoder-based architectures (Yuan et al. 2023), which integrated retrieved text and images with query modalities. With Large Vision-Language Models (LVLMs), (He et al. 2024a) proposed retrieving relevant labeled examples during inference. (Xia et al. 2024b) addressed LVLMs’ misalignment with medical knowledge using retrieval-augmented generation to fetch similar medical reports, employing contrastive learning and context selection strategies. (Xia et al. 2024a) proposed domain-aware retrieval for diverse medical data sources, while (Wu et al. 2025) combined retrieval augmentation with knowledge graphs to reduce hallucination.
Method
Figure 3: Jomed combines a general-purpose VLLM with a general-purpose multi-modal retriever in a two-phase approach. First, with a frozen retriever, we augment training by prepending retrieved (image, caption/report) pairs to original queries, training the LVLM on these augmented prompts. Second, with the LVLM frozen, we sequentially train the dual-head multi-modal retriever using Sharp Perplexity Distillation loss (KL-divergence between the distribution derived from the model logits and the distribution derived from the retrieved candidates’ similarity).
@@
Overview. In this section, we present our methodology, named Jomed (Joint Optimization of Multimodal Retrieval for Enhanced Medical Diagnosis). Our task involves classifying medical images and medical visual question answering. Jomed consists of two main components: a multimodal retriever and a reader. We are given an input patient image and a diagnostic question. The multimodal retriever is tasked with identifying relevant medical knowledge from either literature or hospital records—in the form of images and associated captions or hospital reports that provide predictive information. The reader then analyzes the retrieved candidates along with the diagnostic question and generates an answer. We jointly optimize the reader to better utilize retrieved candidates, and the multimodal retriever to improve its ability to retrieve informative candidates for the reader. Figure 3 illustrates our method. The multimodal retriever has two retrieval heads, textual and visual, for finding candidates in an index of image-caption pairs derived from literature and hospital records. Given an image query, these retrievers find candidates that are then passed to the reader.
@@
Unlike previous work (He et al. 2024a; Xia et al. 2024b, a) our method does not use medical pretraining, and instead we use readily available general LVLMs (Pixtral and Qwen2-vl) and a general retriever, jina-clip (Xiao, Mastrapas, and Wang 2024). The selected LVLMs can process multiple images and texts, allowing us to incorporate both the retrieved image and the retrieved caption/report into the model context. We also compare our results to Med-Flamingo (Moor et al. 2023). MedFlamingo is a medical LVLM which is also able to process multiple images, unlike other medical LVLM available at the time of conducting our experiments. Finally, we also explore the utility of o3 (OpenAI 2025) as a reranker. We now elaborate on the details.
Reader Retrieval Augmentation Fine-Tuning
@@
For an image, question tuple $(i_{d},q_{d})$ $d\in\{1,\ldots,D\}$ from dataset $D$ , we retrieve $K$ relevant tuples of images and texts (reports or captions) $(i_{k},t_{k})$ $k\in\{1,\ldots,K\}$ using our multimodal retriever module. These retrieved pairs are prepended to the image-question pair $(i_{d},q_{d})$ to create augmented inputs. The augmented input consists of the retrieved context followed by the target question-image pair: $(i_{k},t_{k}),(i_{d},q_{d})$ $k\in\{1,\ldots,K\}$. To the best of our knowledge, we are the first to use both the retrieved image and caption/report in context within medical RAG for LVLMs. We use supervised fine-tuning on the augmented input, minimizing the negative log-likelihood to predict the answer $a_{d}$:
$\mathcal{L}(\theta)=-\sum_{d=1}^{D}\sum_{k=1}^{K}\log p_{\theta}(a_{d}|(i_{k},t_{k})\circ(i_{d},q_{d}))$
@@
where $\theta$ represents the parameters of the LVLM, and $\circ$ denotes the concatenation operation in the context window.
Sequential Multimodal Retriever Fine-Tuning
Our dual-head multimodal retriever has a text-based head for retrieving relevant (image, text) pairs based on textual similarity to the query, and an image-based head for retrieving pairs based on visual similarity. To train this multimodal retriever, we measure how each retrieved (image, text) pair contributes to the LVLM’s performance, optimizing the retriever to prioritize high-contribution candidates.
@@
We adopt a sequential training strategy, first optimizing the text retrieval head followed by the image retrieval head. During this process, the reader model remains frozen, focusing solely on improving the retriever’s embedding space. We formulate a Sharp Perplexity Loss as a simple variant of the perplexity distillation loss (Izacard et al. 2023). Similarly to Izacard et al. (2023) we minimize the KL-divergence between the LVLM’s posterior distribution over retrieved pairs and the retriever’s distribution, however we do so directly on downstream tasks with only lightweight fine-tuning in the data-efficient regime without pre-training.
@@
The posterior distribution reflects the model’s confidence in predicting the correct answer $\mathbf{a}$ given a retrieved (image, text) pair $(\mathbf{i}_{k},\mathbf{t}_{k})$ and a query image-question pair $(\mathbf{i},\mathbf{q})$:
$p_{k}\propto p_{\text{LVLM}}(\mathbf{a}\mid(\mathbf{i}_{k},\mathbf{t}_{k})\circ(\mathbf{i},\mathbf{q})),$
@@
where $p_{\text{LVLM}}(\mathbf{a}\mid(\mathbf{i}_{k},\mathbf{t}_{k})\circ(\mathbf{i},\mathbf{q}))$ is the probability assigned by the LVLM to $\mathbf{a}$, and $\circ$ denotes prepending the retrieved chunk to the query.
@@
We sharpen the LVLM’s output distribution by restricting it to only the relevant class tokens from the benchmark’s possible answers. Specifically, we extract the model’s logits, select only the relevant tokens for each class, and apply a softmax to obtain a more discriminative distribution (for VQA tasks, we use yes/no tokens and skip open questions.). We denote this class-restricted distribution as $p_{\text{LVLM}_{C}}$ where $C=\{c_{1},c_{2},c_{3},\ldots\}$ is the set of chosen class tokens. The normalized posterior $p_{k}$ is formulated as:
$p_{k}=\frac{\exp(\log p_{\text{LVLM}_{C}}(\mathbf{a}\mid(\mathbf{i}_{k},\mathbf{t}_{k})\circ(\mathbf{i},\mathbf{q})))}{\sum_{i=1}^{K}\exp(\log p_{\text{LVLM}_{C}}(\mathbf{a}\mid(\mathbf{i}_{i},\mathbf{t}_{i})\circ(\mathbf{i},\mathbf{q})))},$
@@
$p_{\text{RETR}}((\mathbf{i},\mathbf{t})\mid(\mathbf{i},\mathbf{q}))=\frac{\exp(s((\mathbf{i},\mathbf{t}),(\mathbf{i},\mathbf{q}))/\tau)}{\sum_{k=1}^{K}\exp(s((\mathbf{i}_{k},\mathbf{t}_{k}),(\mathbf{i},\mathbf{q}))/\tau)},$
where $s((\mathbf{i},\mathbf{t}),(\mathbf{i},\mathbf{q}))$ is the similarity score between the pair and the query, and temperature $\tau$ controls distribution sharpness. For the text retrieval head, the score is computed using the dot product between the index caption/report embeddings and the query image embedding, while for the image retrieval head, we use the index image embeddings.
Finally, for our loss we compute the KL-divergence between retriever and reader distributions $\text{KL}(p_{\text{LVLM}_{C}}\|p_{\text{RETR}})$:
$\sum_{k=1}^{K}p_{\text{LVLM}_{C}}((\mathbf{i}_{k},\mathbf{t}_{k}))\log\left(\frac{p_{\text{LVLM}_{C}}((\mathbf{i}_{k},\mathbf{t}_{k}))}{p_{\text{RETR}}((\mathbf{i}_{k},\mathbf{t}_{k}))}\right)$
Inference
Following Shi et al. (2023) for a given (image, question) pair, we retrieve the top-$k$ relevant (image, text) chunks using the image as the query. Each chunk is prepended to the question, and the LVLM computes predictions for the augmented prompts in parallel. The final output probability is:
$p_{\text{LVLM}}(a\mid\mathbf{q})=\sum_{k=1}^{K}p_{\text{LVLM}}(a\mid(\mathbf{i}_{k},\mathbf{t}_{k})\circ\mathbf{q})\cdot p_{\text{R}}((\mathbf{i}_{k},\mathbf{t}_{k})\mid\mathbf{q}),$
where $p_{\text{R}}((\mathbf{i}_{k},\mathbf{t}_{k})\mid\mathbf{q})$ is:
$p_{\text{R}}((\mathbf{i}_{k},\mathbf{t}_{k})\mid\mathbf{q})=\frac{\exp(s(\mathbf{q},(\mathbf{i}_{k},\mathbf{t}_{k})))}{\sum_{j=1}^{K}\exp(s(\mathbf{q},(\mathbf{i}_{j},\mathbf{t}_{j})))},$
and $s(\mathbf{q},(\mathbf{i}_{k},\mathbf{t}_{k}))$ is the similarity score between the query and the retrieved chunk.
Backbone Model Breast Derma Retina VinDr-PCXR BRSET Mean

### v3

# Lightweight Joint Optimization of General-Purpose Vision-Language Models and Retrievers for RAG-Based Medical Diagnosis

@@ Abstract
Retrieving relevant visual and textual information from medical literature and hospital records can enhance diagnostic accuracy for clinical image interpretation. We develop a multimodal retrieval model jointly optimized with an LVLM for medical diagnosis, unlike standard RAG which doesn’t backpropagate LVLM errors to the retriever. Using only general-purpose backbones with lightweight fine-tuning, our model achieves competitive results with medically-pretrained models on clinical classification and VQA tasks. In a novel analysis, we find that different top-retrieved images often yield different predictions for the same target, and that these cases are challenging for all models, even for non-retrieval models. Our joint retrieval optimization significantly improves these cases over standard RAG. However, oracle analysis reveals that while the correct diagnosis is frequently achievable using one of the top retrieved images, in practice there is a large performance gap from the oracle, and rerankers using frontier LVLMs do not close this gap—leaving ample room for improvement by future methods. Code available at https://github.com/Nirmaz/JOMED.
Lightweight Joint Optimization of General-Purpose Vision-Language Models and Retrievers for RAG-Based Medical Diagnosis
Nir Mazor1, Tom Hope1,2 1School of Computer Science and Engineering, The Hebrew University of Jerusalem 2The Allen Institute for AI (AI2)
1 Introduction
@@
Figure 1: We jointly optimize a multimodal retriever and a Large Vision-Language Model for medical tasks. We achieve competitive results without resource-intensive medical pre-training and significantly improve performance on challenging cases where different retrieved images lead to inconsistent retrieval prediction.
@@
Inferring diagnoses from medical imagery is a fundamental part of clinical decision-making. Large Vision Language Models (LVLMs) have been widely explored for medical diagnosis Thawakar et al. (2024); Wu et al. (2023); Zhang et al. (2023b); Moor et al. (2023); Li et al. (2023). To improve the performance of LVLMs in the medical field, retrieval augmentation (RAG) has been adopted and has shown promising results, providing more accurate and also explainable methods He et al. (2024a); Xia et al. (2024b, a); Wu et al. (2025).
@@
In this work, we explore a lightweight fine-tuning approach for taking a general-purpose LVLM and optimizing it together with a general-purpose multimodal retriever for medical diagnosis tasks. Different from the common multimodal RAG setup, we perform joint optimization of the LVLM and retriever. This teaches the LVLM to utilize retrieved knowledge—images, captions and clinical reports from both literature and hospital records—while at the same time tuning the retriever to find information that leads to correct predictions by the LVLM. Poorly optimized retrieval mechanisms can mislead models Yoran et al. (2023); Sun et al. (2024); as shown in Figure 2, a RAG model incorrectly diagnoses a benign ultrasound when using a cancerous retrieved image, but joint optimization leads to correct retrieval and prediction.
@@
Our method involves sequential multimodal training with a dual-head retriever architecture and a customized retrieval loss and visual question answering (VQA) training recipe. Our method achieves competitive results in medical image classification and VQA. Importantly, our focus in this work is not to chase SOTA results but to shed light on several interesting observations. One, that a simple and effective joint LVLM-retriever optimization mechanism can lead to a substantial boost in results over current multimodal RAG methods and should thus be more widely adopted for medical diagnosis. Second, that this mechanism achieves competitive results by using models with no medical pre-training (for neither the LVLM nor the retriever) and only lightweight fine-tuning. This resonates with recent findings showing general-purpose LVLMs can rival their medical counterparts Jeong et al. (2024) and has potentially important implications considering the resource-intensive nature of pre-training processes.
@@
Third, previous work on joint optimization in the general domain required large-scale pre-training followed by few-shot task-specific adaptation Izacard et al. (2023); Hu et al. (2023); Lin et al. (2024). In contrast, we conduct joint optimization directly on downstream tasks with no pretraining and only lightweight fine-tuning (as few as 546 samples). This provides first evidence about the feasibility and utility of lightweight data-efficient joint optimization directly on downstream tasks as opposed to in the pre-training setting. We are also the first to explore and show the utility of joint optimization in the multimodal medical domain.
@@
Fourth and importantly, we conduct a novel analysis that finds that our method helps in particular to address a class of errors we term inconsistent retrieval predictions. Inconsistent retrieval predictions are cases in which for a given patient image, each retrieved image leads the model to make different predictions. These cases commonly occur across our experiments and are substantially more difficult for models, both retrieval-augmented and non-retrieval models. This instability with respect to different retrieval candidates leads to high prediction entropy/uncertainty across classes, degrading not only overall RAG results but also their reliability Lambert et al. (2022). As we show, our joint optimization mechanism significantly mitigates this issue and achieves large improvement over standard fine-tuned RAG on these cases. We further discover that for a large proportion of these inconsistent cases, at least one retrieved candidate enables the model to predict the correct answer in an oracle setting in which we provide the model with the correct retrieved candidate. This suggests useful information is being retrieved but is often lost among less helpful candidates. To address this, we explored using the powerful multimodal model o3 OpenAI (2025) as a reranker to identify the most predictive instances. We found it did not close the performance gap to the oracle and was comparable or inferior to our method, leaving significant room for future improvement.
@@
  • We jointly optimize a multimodal retriever and an LVLM for medical classification and VQA. We use only general-purpose backbones without any medical pretraining, and show that by retrieving medical information and co-optimizing the retriever and LVLM we can close much of the gap to expensive medically pre-trained models.
@@
  • Unlike previous general-domain joint optimization that required pre-training, we conduct joint optimization directly on downstream tasks with only lightweight data-efficient fine-tuning.
@@
  • We identify inconsistent retrieval predictions, challenging cases where different retrieved images lead to different predictions for the same query. Our joint optimization method significantly improves performance on these cases when compared to standard RAG.
@@
Figure 2: joint optimization impacts inconsistent retrieval predictions. After joint optimization, retrieved images show greater alignment with query image labels. In Breast, a cancer-free ultrasound query (a) initially retrieved a lesion image (b), however, after joint optimization, the retrieved image is less directly related to the wrong label (c). In VQA-RAD, retrieval shifted from an unrelated medical condition (e) to an image depicting the condition of the query image.
2 Related Work
@@
Retrieval Augmented joint optimization has been explored primarily in unimodal (text only) NLP with work such as ATLAS Izacard et al. (2023), which conducted large-scale pretraining of both reader and retriever models and evaluated in zero-shot and few-shot scenarios in the general domain. REVEAL Hu et al. (2023) extended this to the multimodal setting, achieving state-of-the-art results in visual tasks in the general domain with extensive pre-training. Shi et al. (2023) proposed training textual retrievers based on reader performance, influencing subsequent joint-optimization methods. Lin et al. (2024) introduced another joint optimization variant and proposed sequential training: first training the reader with retrieval augmentation, then training the retriever. Siriwardhana et al. (2023) demonstrated joint optimization’s effectiveness for domain adaptation. Our work differs by being the first to explore joint optimization optimized directly for downstream tasks, with minimal fine-tuning, as opposed to requiring large-scale pre-training. Our extensive novel analyses are the first to shed light on inconsistent predictions and how joint optimization on downstream tasks helps address them.
@@
Multimodal Retrieval Augmentation in Medical Applications. Multimodal retrieval augmentation began with encoder-based architectures Yuan et al. (2023), which integrated retrieved text and images with query modalities. With Large Vision-Language Models (LVLMs), He et al. (2024a) proposed retrieving relevant labeled examples during inference. Xia et al. (2024b) addressed LVLMs’ misalignment with medical knowledge using retrieval-augmented generation to fetch similar medical reports, employing contrastive learning and context selection strategies. Xia et al. (2024a) proposed domain-aware retrieval for diverse medical data sources, while Wu et al. (2025) combined retrieval augmentation with knowledge graphs.
3 Method
Figure 3: The two-phase Jomed training. The LVLM is first trained with a frozen retriever on augmented prompts containing retrieved image–caption/report pairs. With the LVLM frozen, the dual-head multimodal retriever is optimized using a KL divergence computed over a selected subset of the model’s logits. An LLM converts open-ended questions into closed form (during training only), which improves results.
@@ 3 Method
Overview. In this section, we present our methodology, termed Jomed (Joint Optimization of Multimodal Retrieval for Enhanced Medical Diagnosis). Our task involves medical image classification and visual question answering. Jomed comprises two main components: a multimodal retriever and a reader. Given an input patient image and a diagnostic question, the multimodal retriever identifies relevant medical knowledge—images and their associated captions or hospital reports—that provide predictive information. The reader then analyzes the retrieved candidates along with the question to generate an answer. We jointly optimize the reader to better leverage the retrieved information and the multimodal retriever to enhance its ability to select informative candidates for the reader. Figure 3 illustrates our framework.
@@
Unlike previous work He et al. (2024a); Xia et al. (2024b, a) our method does not use medical pretraining, and instead we use readily available general LVLMs (Pixtral and Qwen2-vl) and a general retriever, jina-clip Xiao et al. (2024). The selected LVLMs can process multiple images and texts, allowing us to incorporate both the retrieved image and the retrieved caption/report into the model context. We also compare our results to Med-Flamingo Moor et al. (2023). Med-Flamingo is a medical LVLM which is also able to process multiple images, unlike other medical LVLM available at the time of conducting our experiments. We now elaborate on the details.
3.1 Reader Retrieval Augmentation Fine-Tuning
@@
For an image-question tuple $(i_{d},q_{d})$ where $d\in\{1,\ldots,D\}$ from dataset $D$, we retrieve $K$ relevant tuples of images and texts (reports or captions) $(i_{k},t_{k})$ where $k\in\{1,\ldots,K\}$ using our multimodal retriever module. To simplify notation, we denote $z_{k}=(i_{k},t_{k})$ as the $k$-th retrieved image-text pair and $q_{d}=(i_{d},q_{d})$ as the query pair for example $d$. These retrieved pairs are prepended to the query pair to create augmented inputs. The augmented input consists of the retrieved context followed by the target question-image pair: $z_{k}\circ q_{d}$ for $k\in\{1,\ldots,K\}$. We use supervised fine-tuning on the augmented input, minimizing the negative log-likelihood to predict the answer $a_{d}$:
$\mathcal{L}(\theta)=-\sum_{d=1}^{D}\sum_{k=1}^{K}\log p_{\theta}(a_{d}\mid z_{k}\circ q_{d})$
@@
where $\theta$ represents the parameters of the LVLM, and $\circ$ denotes the concatenation operation.
3.2 Sequential Multimodal Retriever Fine-Tuning
@@
Our dual-head multimodal retriever has a text-based head for retrieving relevant (image, text) pairs based on textual similarity to the query, and an image-based head for retrieving pairs based on visual similarity. To train this retriever, we adopt a sequential training strategy, first optimizing the text retrieval head followed by the image retrieval head. During this process, the reader model remains frozen, focusing solely on improving the retriever’s embedding space. Similarly to Izacard et al. (2023) we minimize the KL divergence between the LVLM’s posterior distribution over retrieved pairs and the retriever’s distribution, however we do so directly on downstream tasks with only lightweight fine-tuning in the data-efficient regime without pre-training, and we customize the loss and the training recipe for open-ended questions leading to empirical gains.
@@
Let $z_{k}=(\mathbf{i}_{k},\mathbf{t}_{k})$ denote the $k$-th retrieved image-text pair and $q=(\mathbf{i},\mathbf{q})$ denote the query pair. The posterior distribution reflects the model’s confidence in predicting the correct answer $\mathbf{a}$ given a retrieved pair $z_{k}$ and query $q$:
$p_{k}\propto p_{\text{LVLM}}(\mathbf{a}\mid z_{k}\circ q),$
@@
where $p_{\text{LVLM}}(\mathbf{a}\mid z_{k}\circ q)$ is the probability assigned by the LVLM to $\mathbf{a}$, and $\circ$ denotes prepending the retrieved chunk to the query.
@@
We sharpen the LVLM’s output distribution by restricting it to only the relevant class tokens from the benchmark’s possible answers. Specifically, we extract the model’s logits, select only the relevant tokens for each class, and apply a softmax to obtain a more discriminative distribution. We denote this class-restricted distribution as $p_{\text{LVLM}_{C}}$ where $C=\{c_{1},c_{2},c_{3},\ldots\}$ is the set of chosen class tokens. The normalized posterior $p_{k}$ is formulated as:
$p_{k}=\frac{\exp(\log p_{\text{LVLM}_{C}}(\mathbf{a}\mid z_{k}\circ q))}{\sum_{i=1}^{K}\exp(\log p_{\text{LVLM}_{C}}(\mathbf{a}\mid z_{i}\circ q))},$
@@
$p_{\text{RETR}}(z\mid q)=\frac{\exp(s(z,q)/\tau)}{\sum_{k=1}^{K}\exp(s(z_{k},q)/\tau)},$
where $s(z,q)$ is the similarity score between the pair and the query, and temperature $\tau$ controls distribution sharpness. For the text retrieval head, the score is computed using the dot product between the index caption/report embeddings and the query image embedding, while for the image retrieval head, we use the index image embeddings.
For our loss we compute the KL-divergence between retriever and reader $\text{KL}(p_{\text{LVLM}_{C}}\|p_{\text{RETR}})$:
$\sum_{k=1}^{K}p_{\text{LVLM}_{C}}(z_{k})\log\left(\frac{p_{\text{LVLM}_{C}}(z_{k})}{p_{\text{RETR}}(z_{k})}\right)$
Finally, for visual question answering (VQA), we use the o3 model to convert open-ended questions into closed-ended ones (yes/no). See prompt in Appendix E and analysis of approach contribution in Appendix D.1. We apply this only during training [^1] Run once per question; the total cost is only a few dollars., without modifying questions at inference time. Reformatted questions will be made publicly available. This reformulation improves results, akin to restricting the LVLM output distribution.
3.3 Inference
Following Shi et al. (2023) for a given query pair $q=(\mathbf{i},\mathbf{q})$, we retrieve the top-$K$ relevant chunks $z_{k}=(\mathbf{i}_{k},\mathbf{t}_{k})$ using the image as the query. Each chunk is prepended to the question, and the LVLM computes predictions for the augmented prompts in parallel. The final output probability is:
$p_{\text{LVLM}}(a\mid q)=\sum_{k=1}^{K}p_{\text{LVLM}}(a\mid z_{k}\circ q)\cdot p_{\text{R}}(z_{k}\mid q),$
where $p_{\text{R}}(z_{k}\mid q)$ is:
$p_{\text{R}}(z_{k}\mid q)=\frac{\exp(s(q,z_{k}))}{\sum_{j=1}^{K}\exp(s(q,z_{j}))},$
and $s(q,z_{k})$ is the similarity score between the query and the retrieved chunk.
Backbone (Size) Model Breast Derma Retina VinDr-PCXR BRSET Mean